Homework 1 - Comparative Advantage st. Length 2:0000 Shivamkumar Patel: Altempt 1 Both France and Spain Neither France nor Spain Page 1: Question 14 (1 point) Suppose that Spain and France both produce ships and grapes, both countries. The table below shows the combinations of the two goods that each country can produce in one year using the same amounts of capital and labor. Suppose that Spain is currently producing 30 ships and 30 tons of grapes, and France is currently producing 30 ships and 30 tons of which are sold for the same price in If Spais and France each specialsze in producing only one good (the good for which each has a ave advantage), then how many additional tons of grapes can be produced for the two countries 10 1112 combined? Provide your answer as a namber rounded to two decirsal places. Use amounts Do not mclude any symbols, such as S%in your answer to indicate segative 13 141 Grapes (in tons) 50 40 30 20 10 16 17 75 10 20 30 45 30 15 15 30 19 20 2 75 22 23 Your Answer
Answer #1

Before specialization:

Total production of grapes = 60 tons

Total production of ships = 60

After specialization:

Spain - opportunity cost of producing 1 ship = 75/50 = 1.5 tons of grapes

France - opportunity cost of producing 1 ship = 50/75 = 0.66 ton of grapes

France has a lower opportunity cost of producing 1 ship and it would specialize in Ship production (it has a comparative advantage in ship production)

Total production of ships = 75

Total production of grapes = 75 tons

Additional production of grapes = 75 - 60 = 15 tons